Sally is Bob's younger sister, and she is 7 years younger than Bob. If you multiply Bob's age by three, and subtract sally's age, you get their mother's age, which is 37. How old is Sally?
Answer by TheInstructor(29)

Let Bob = x

Sally = x-7

Multiply Bob's age by 3: 3x
Subtract Sally's age: 3x-(x-7) = 2x+7

2x+7 = 37
Therefore x= 15
Bob = 15
Sally = 8

Hope that helped! Feel free to leave a message for any inquiries! =]